### v1

@@ 4 Revisit on The Fidelity Term of (5)
    $G^{*}=\Omega_{1}\cup\Omega_{2}$, when (21) is satisfied.
5 Method
In this section, we describe our proposed Quasi-conformal Interactive Segmentation model in detail.
5.1 Overall Description
Figure 3: A flow chart illustrating the process of the quasiconformal interactive segmentation (QIS) model.
In this subsection, we provide an overview of the interactive segmentation procedure. Figure 3 visually illustrates the process. Initially, the Quasi-conformal registration-based segmentation model, described in (6), is applied to obtain an initial segmentation. This initial segmentation serves as the current segmentation mask, which may not be entirely accurate and thus requires further adjustments. To guide these adjustments, the user can indicate regions to be included (using positive clicks) and regions to be excluded (using negative clicks). These clicks are then converted into a click map, which is subsequently incorporated into the Quasi-conformal segmentation model. By minimizing the interactive segmentation energy, a segmentation Quasi-conformal map is generated, resulting in an improved segmentation mask. If the improved segmentation is not satisfactory, the process is repeated until a satisfactory segmentation result is achieved. This iterative procedure, which incorporates the user’s input, is referred to as the interactive segmentation process.
In the following subsections, we will discuss the definitions of positive clicks, negative clicks, and click maps in details. Additionally, we will provide a comprehensive explanation of the proposed interactive segmentation fidelity term. This fidelity term is a crucial component that is integrated into the Quasi-conformal segmentation model, specifically designed for interactive segmentation purposes.
5.2 Click map
In this subsection, we introduce the click map, which is essential for constructing our proposed interactive segmentation model.
In this work, we propose a methodology to generate the click map from user clicks based on $K$-means clustering. Suppose a collection $\mathcal{C}_{*}$ of positive clicks or negative clicks is given. In other words, $\mathcal{C}_{*}=\mathcal{C}_{p}$ or $\mathcal{C}_{n}$. The image domain $\Omega$ can be decomposed into $K$ clusters based on the image intensity $I:\Omega\to\mathbb{R}$, using the $K$-means clustering. Mathematically, we divide $\Omega$ into $K$ sets, namely, $\mathcal{K}_{1}$, $\mathcal{K}_{2}$, $\dots$, $\mathcal{K}_{K}$ by mininizing the intra-cluster variance:
$\min_{\mathcal{K}_{1},\cdots,\mathcal{K}_{K}}\sum_{i=1}^{K}\sum_{\bm{x}\in\mathcal{K}_{i}}\|I(\bm{x})-I(\bm{\mu}_{i})\|^{2},$ (22)
where $\bm{\mu}_{i}$ is the centroid of the cluster $i$. This expression quantifies the similarity of data points within each cluster. The $K$-means clustering can be computed using the following algorithm.
Algorithm 1 K-means Clustering Segmentation
  Choose the number of clusters $K$;
  Randomly select $K$ points from $\Omega$ as $\bm{\mu}_{1}$$\dots$,$\bm{\mu}_{K}$;
  for $l=1:\text{MaxKS}$ do
     Solve (22) to obtain $\mathcal{K}_{1}$, $\cdots$, $\mathcal{K}_{K}$;
     Recompute $\bm{\mu}_{i}$ to be the centroid of $\mathcal{K}_{i}$ for each $i$;
     if stopping criteria is satisfied then
        berak;
     end if
  end for
Note that each cluster $\mathcal{K}_{i}$, may consist of multiple isolated components. We denote the isolated components by $D^{i}_{j}$. Hence, $\mathcal{K}_{i}=\bigcup_{j=1}^{N_{i}}D^{i}_{j}$, where $N_{i}$ indicates the number of isolated regions within cluster $i$. For each $\bm{x}_{\text{click}}\in\mathcal{C}_{*}$, we define $\delta(\bm{x}_{click})$ as follows:
$\delta(\bm{x}_{click})={D^{i}_{j}}\text{ for }(i,j)\text{ if }\bm{x}_{click}\in D^{i}_{j}.$ (23)
A local homogeneous region surrounding clicks can be obtained by taking the union, which is given by $\mathcal{H}=\bigcup_{\bm{x}_{click}\in\mathcal{C}_{*}}\delta(\bm{x}_{click})$. A click map $M_{\text{click}}$, which is a binary map capturing $\mathcal{H}$, can be defined as follows.
Definition 5.1 (Click map).

### v2

@@ 4.3.1 Revisit on the fidelity term of (3)
    when (14) is satisfied, $G^{*}=\Omega_{1}\cup\Omega_{2}$.
4.3.2 User’s Clicks and Parameter Selection
Given a click map $M_{\mathrm{clicks}}:\Omega\longrightarrow\mathbb{R}$, our proposed interactive segmentation fidelity energy is defined as
(15) $E_{int}(c_{1},c_{2},G;M_{\mathrm{clicks}})=\int_{\Omega}(I+rM_{\mathrm{clicks}}-c_{1}\mathcal{X}_{G}-c_{2}\mathcal{X}_{G^{c}})^{2}\mathrm{d}\bm{x}.$
Case 1: Positive click.
Mathematically, suppose the region of interest is represented by the union of two regions, $\Omega_{1}\cup\Omega_{2}$. However, the initial segmentation only identifies $\Omega_{1}$, omitting the presence of $\Omega_{2}$. To rectify this, a positive click is required to indicate that $\Omega_{2}$ should also be included as part of the region of interest. Therefore, by considering the click map $M_{\mathrm{clicks}}=\mathcal{X}_{\Omega_{2}}$ and leveraging the interactive segmentation fidelity energy, the parameter $r$ in (15) needs to be carefully selected to ensure that the updated model can generate an optimal segmentation $G^{**}=\Omega_{1}\cup\Omega_{2}$. The following theorem describes how $r$ can be chosen.
Theorem 4.5.
Given a three-value image $I$ defined as (8), with the initial segmentation $G^{*}=\Omega_{1}$ derived by (7) and the click map $M_{\mathrm{clicks}}=\mathcal{X}_{\Omega_{2}}$, we can obtain the desired segmentation $G^{**}=\Omega_{1}\cup\Omega_{2}$ from the interactive segmentation fidelity energy (15) if $r$ satisfies
(16) $\begin{cases}\frac{(p_{1}-p_{2})B-(p_{0}-p_{1})}{B}\leq r\leq\frac{(p_{1}-p_{2})A+(p_{0}-p_{2})}{A+1}\ \text{when $p_{0}>p_{1}$},\\ \frac{(p_{1}-p_{2})A+(p_{0}-p_{2})}{A+1}\leq r\leq\frac{(p_{1}-p_{2})B-(p_{0}-p_{1})}{B}\ \text{when $p_{0}<p_{1}$},\end{cases}$
where $A=\sqrt{\frac{||\Omega_{1}||(||\Omega_{2}||+||\Omega_{0}||)}{||\Omega_{0}||(||\Omega_{1}||+||\Omega_{2}||)}}$ and $B=\sqrt{\frac{||\Omega_{2}||(||\Omega_{1}||+||\Omega_{0}||)}{||\Omega_{0}||(||\Omega_{1}||+||\Omega_{2}||)}}$. In practice, $r$ can be chosen as the middle point of the range, i.e. $r=\left(\frac{(p_{1}-p_{2})A+(p_{0}-p_{2})}{A+1}+\frac{(p_{1}-p_{2})B-(p_{0}-p_{1})}{B}\right)/2$.
Proof 4.6.
See Appendix B.
Case 2: Negative click.
When the initial segmentation generated by (7) includes an additional region that is not part of the desired segmentation, a negative click is required to remove it. Assume that the region of interest is denoted as $\Omega_{1}$, but the initial segmentation erroneously includes an extra region $\Omega_{2}$. By considering the click map $M_{\mathrm{clicks}}=\mathcal{X}_{\Omega_{2}}$, we need to carefully choose the parameter $r$ in (15) to ensure that the updated model can produce an optimal segmentation $G^{**}=\Omega_{1}$. The following theorem provides details on how the value of $r$ can be selected.
Theorem 4.7.
Given a three-value image $I$ defined as (8), with the initial segmentation $G^{*}=\Omega_{1}\cup\Omega_{2}$ derived from (7) and the click map $M_{\mathrm{clicks}}=\mathcal{X}_{\Omega_{2}}$, we can obtain the desired segmentation $G^{**}=\Omega_{1}$ from the interactive segmentation fidelity energy (15) if $r$ satisfies
(17) $\begin{cases}\frac{(p_{0}-p_{2})C+(p_{1}-p_{2})}{C+1}\leq r\leq\frac{(p_{0}-p_{2})E+(p_{0}-p_{1})}{E}\ \text{when $p_{0}>p_{1}$},\\ \frac{(p_{0}-p_{2})E+(p_{0}-p_{1})}{E}\leq r\leq\frac{(p_{0}-p_{2})C+(p_{1}-p_{2})}{C+1}\ \text{when $p_{0}<p_{1}$},\end{cases}$
where $C=\sqrt{\frac{||\Omega_{0}||(||\Omega_{2}||+||\Omega_{1}||)}{||\Omega_{1}||(||\Omega_{0}||+||\Omega_{2}||)}}$ and $E=\sqrt{\frac{||\Omega_{2}||(||\Omega_{0}||+||\Omega_{1}||)}{||\Omega_{1}||(||\Omega_{0}||+||\Omega_{2}||)}}$. In practice, $r$ can be chosen as the middle point of the range, i.e. $r=\left(\frac{(p_{0}-p_{2})C+(p_{1}-p_{2})}{C+1}+\frac{(p_{0}-p_{2})E+(p_{0}-p_{1})}{E}\right)/2$
Proof 4.8.
See Appendix C.
Our method is an interactive model that involves human input for each step, which leads to some factors in the whole procedure being undetermined. However, in each step where the clicks are given to be negative or positive, the click map can then be obtained, and so is the parameter $r$. With those components fixed, in each step $n$ of the proposed interactive segmentation model, the following theorem can guarantee the existence of the minimizer.
Theorem 4.9 (Existence of minimizer).
For each step $n$ of the proposed interactive segmentation model, suppose $\Omega$ is bounded and simply connected, $I^{n-1}$ is a function from $\Omega\subset\mathbb{R}^{n}\rightarrow\mathbb{R}$, $r^{n}$ is chosen by (16) or (17), $M^{n}_{\mathrm{clicks}}$ is computed by (5) and $\alpha_{i}>0,i=1,2$. Let
$\mathcal{A}:=\{\varphi\in\mathcal{C}^{2}(\Omega):\|\varphi\|_{\infty}\leq d_{1},\|\nabla\varphi\|_{\infty}\leq d_{2},\|\nabla\varphi^{2}\|_{\infty}\leq d_{3},\det\nabla\varphi\geq\epsilon\}$
for some $d_{i}>0,i=1,2,3$ and a small $\epsilon>0$. Then each step $n$ of the proposed interactive segmentation model admits a minimizer in $\mathcal{A}$. In fact, $\mathcal{A}$ is compact.
Proof 4.10.
The proof is similar with Theorem 4.1 in [39] and [47].
5 Numerical implementation
In this section, we give the details of how to solve the proposed interactive model in Section 4.1.
The main components in each step $n$ of the proposed model are to compute the click map $M^{n}_{\mathrm{clicks}}$, choose the weight $r^{n}$ and solve the following variational problem
(18) $\min_{c_{1},c_{2},\varphi}\frac{1}{2}\int_{\Omega}(I^{n}-c_{1}\mathcal{X}_{D}\circ\varphi-c_{2}\mathcal{X}_{D^{c}}\circ\varphi)^{2}\mathrm{d}\bm{x}+\alpha_{1}\int_{\Omega}|\Delta\varphi|^{2}\mathrm{d}\bm{x}+\alpha_{2}\int_{\Omega}\phi(|\mu(\varphi)|^{2})\mathrm{d}\bm{x},$
where $I^{n}=I^{n-1}+r^{n}M^{n}_{\mathrm{clicks}}$. In Section 4.2 and 4.3, we have provided the approach to calculate the click map $M^{n}_{\mathrm{clicks}}$ and the corresponding $r^{n}$, respectively. The scaling function that used to better penalize large $v=|\mu(\varphi)|$ is chosen as $\phi(v)=\frac{1}{(v-1)^{2}}$ [46]. Next, we mainly investigate the numerical solver for (18).